OpenStudy (anonymous):

How do you get the standard form by using the points (50,475) (10,150)? Find the slope, point slope, standard, show work.

OpenStudy (anonymous):

slope is difference in y's/difference in x's = m m= (475-150)/(50-10) = 325/40 = 65/8 point slope: (y-475)=(65/8)*(x-50) standard: 8y-65x=550 (all fractions to integer coefficients)